OpenStudy (anonymous):

3t – 6t + 8 = 7t – 2

OpenStudy (anonymous):

Simplify the equation.

OpenStudy (anonymous):

what is "t"

OpenStudy (anonymous):

Oh. You want us to give you the answer. $$3t - 6t + 8 = 7t - 2\\ -3t +8 = 7t - 2\\ 10 = 10t\\ t = 1$$

OpenStudy (anonymous):

thanks